OpenStudy (anonymous):

Help please Variable y varies directly with x2, and y = 75 when x = 5. What is the value of y when x = 3?

OpenStudy (anonymous):

If y varies directly with \(x^2\), then you can say: \(y = kx^2\), \(k\) is proportionality constant..

OpenStudy (anonymous):

\(y = kx^2\), Now, \(y = 75\) when \(x = 5\), by using these details, find \(k\) first.

OpenStudy (anonymous):

\(75 = k (5)^2\) \(75 = k \times 25\) Can you find \(k\) here? If so, then tell me.